OpenStudy (anonymous):

Solving by Factoring 3x^2 = -9x What are the solutions for x? x = 3, x = 9 x = 0, x = 3 x = 0, x = -3 none of the above

OpenStudy (anonymous):

3x^2+9x=0 x(3x+9)=0 x=0 ; 3x+9=0 x=0 ; 3x=-9 x=0;-3